Use any appropriate method in finding the indicated solution of the following differentialย 

equations.ย 

[2๐‘ฅ๐‘ฆ cos(๐‘ฅ2) โˆ’ 2๐‘ฅ๐‘ฆ + 1]๐‘‘๐‘ฅ + [๐‘ ๐‘–๐‘›(๐‘ฅ2) โˆ’ ๐‘ฅ2]๐‘‘๐‘ฆ = 0


1
Expert's answer
2022-03-31T05:39:40-0400

"\left[ 2xy\cos x^2-2xy+1 \right] dx+\left[ \sin x^2-x^2 \right] dy=0\\\frac{\partial}{\partial x}\left[ y\sin x^2-x^2y+x \right] dx+\frac{\partial}{\partial y}\left[ y\sin x^2-x^2y+x \right] dy=0\\d\left[ y\sin x^2-x^2y+x \right] =0\\y\sin x^2-x^2y+x=C\\y=\frac{C-x}{\sin x^2-x^2}"